OpenStudy (anonymous):

Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 3:35 to 3:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 3π inches? Part 4: What is the coordinate point associated with this radian measure?

OpenStudy (anonymous):

I found 2π/3 for part 1

OpenStudy (anonymous):

8π/3 for part 2

OpenStudy (anonymous):

I also found 3π/4 for part 3, I just need help with part 4

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@beccaboo333

OpenStudy (beccaboo333):

I'm not very good with math ;-;

OpenStudy (anonymous):

Ok sorry

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Would you explain in your own words what you are being asked for in Part 4? What mathematical terms apply?

OpenStudy (anonymous):

I am supposed to use the radian measure that I found in part 3 to find the coordinates of where the minute hand is supposed to be at that measure, but I'm not sure how.

OpenStudy (mathmale):

Good. We need to use the circumference formula. You familiar with that?

OpenStudy (anonymous):

Yes I am. C=πd

OpenStudy (mathmale):

Good. What about the arc length formula? Type that out, please.

OpenStudy (mathmale):

Excuse me, but if you've done Part 3 already, you probably already have some of the answer for Part 4. What is your answer for part 3, and how did y ou obtain it?

OpenStudy (mathmale):

Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 3π inches?

OpenStudy (anonymous):

Arc length is s=r(theta) I worked backwards from the arc length formula

OpenStudy (anonymous):

I did 3π=4(theta)

OpenStudy (anonymous):

And solved for theta

OpenStudy (mathmale):

The tip of the minute had moves 3pi radians along the circle of radius 4 inches. Fair enough: I agree with your logic.

OpenStudy (mathmale):

What is your value for this angle, theta? Are you measuring it in degrees or in radians?

OpenStudy (anonymous):

I got 3π/4 radians, which is 135 degrees.

OpenStudy (anonymous):

It asks for radians in part 3.

OpenStudy (mathmale):

Looks great. So now, the minute hand has rotated (3/4)pi radians. The length of the minute hand is 4 inches. Do these formulas look at all familiar to you? x=r cos theta y=r sin theta

OpenStudy (mathmale):

As a footnote, the cosine function is defined as the length of the side adjacent to the angle in a right triangle, divided by the hypotenuse of the triangle. Familiar?

OpenStudy (anonymous):

No, I'm sorry

OpenStudy (anonymous):

The second part is familiar.

OpenStudy (anonymous):

The footnote, I mean

OpenStudy (mathmale):

Have you studied trigonometry yet? Trig presents the easiest method of coming up with the answer to Part IV.

OpenStudy (anonymous):

Yes I have

OpenStudy (mathmale):

then you know that \[\sin \theta=\frac{ opp }{ hyp }\] where "opp" represents the length of side opposite the angle in question, and "hyp" is (obviously) the hypotenuse length.

OpenStudy (anonymous):

Yes I do

OpenStudy (mathmale):

If\[\sin \theta=\frac{ opp }{ hyp }\]then opp=(hyp)(sin theta).

OpenStudy (mathmale):

\[y=r*\sin \]

OpenStudy (mathmale):

sorry. I mean: y = r * sin theta.

OpenStudy (mathmale):

How do you define the cosine function?

OpenStudy (anonymous):

Do you mean cos=adj/hyp or f(x)=a cos(bx-c)+d ?

OpenStudy (mathmale):

The first formula, that's the definition of the cosine function. If y ou accept that definition, and if you multiply both sides by "hyp," you obtain x= (hyp)(cos theta). Here, your radius is 4 inches, your angle is (3/4)pi or 135 degrees. First find the x-coordinate of the tip of the minute hand by finding x=(4 inches)(cosine 135 deg).

OpenStudy (anonymous):

Is x the x-coordinate that I'm looking for?

OpenStudy (mathmale):

Yes. You want the coordinates of the tip of the minute hand, and are now finding the x-coord. separately. Next, you'll find the y-coord.

OpenStudy (anonymous):

Would the x-coordinate be -2.8?

OpenStudy (mathmale):

Yes! Now, try finding the y-coordinate. Would you expect it to be positive or negative? Why? Hint: in which direction does the minute hand turn?

OpenStudy (anonymous):

It goes clockwise so the y-coordinate would be negative.

OpenStudy (mathmale):

You're doing very well indeed. Indeed the tip of the minute hand would be in quadrant III as measured counterclockwise from 0 degrees or 0 radians. What do you now think are the coordinates of the tip of the minute hand?

OpenStudy (anonymous):

I'm sorry, sir. I'm not sure about what to do from here.

OpenStudy (mathmale):

You've correctly found the x-coordinate of the tip of the minute hand: x=-2.8. You've correctly predicted that the y-coord. is negative. By evaluating (4 inches)(sin [-(3/4)pi], you'll end up with the y-coordinate. What is it?

OpenStudy (anonymous):

I got -0.16 but I don't think that I'm right.

OpenStudy (mathmale):

How did you obtain that? Are you using a calculator?

OpenStudy (anonymous):

Yes I am

OpenStudy (anonymous):

I used sin for 3/4, then multiplied it by pi, and then multiplied it by 4.

OpenStudy (mathmale):

Can you set it to either degree or radian mode? if so, set it to degree mode if you're finding y=sin (-135 deg) or to radian mode if you're finding y = sin ([3/4]pi).

OpenStudy (mathmale):

I see. That explains what happened here. FIRST, multiply (-3/4) by pi. only AFTER that, find the sine of the result.

OpenStudy (mathmale):

Please try again.

OpenStudy (mathmale):

In which mode (degrees or radians) is your calculator currently operating?

OpenStudy (anonymous):

Degree

OpenStudy (mathmale):

then the angle in question is -135 degrees. type in cos (-135) and see what you get.

OpenStudy (anonymous):

I got -0.707

OpenStudy (mathmale):

And that's correct.

OpenStudy (anonymous):

For the y-coordinate?

OpenStudy (mathmale):

Your -0.707 is correct as the value of the sine of (-135 deg). But you still have to multiply that by 4 inches, don't you? Try it, please. Hint: the x- and the y- coordinates happen to be the same here!

OpenStudy (anonymous):

Oh wow, I kept getting -2.8 for y but I thought I was wrong

OpenStudy (anonymous):

So the coordinates are (-2.8.-2.8)?

OpenStudy (mathmale):

Yes, that's perfect, EXCEPT you'll need to write "inches" after each of the two coordinates.

OpenStudy (anonymous):

Ok. Thank you, sir, so very much!

OpenStudy (mathmale):

My great pleasure, fawkes! Talk with you again!